### v1

# Change-Agent: Towards Interactive Comprehensive Change Interpretation and Analysis from Change Detection and Change Captioning

Abstract
Monitoring changes in the Earth’s surface is crucial for understanding natural processes and human impacts, necessitating precise and comprehensive interpretation methodologies. Remote sensing satellite imagery offers a unique perspective for monitoring these changes, leading to the emergence of remote sensing image change interpretation (RSICI) as a significant research focus. Current RSICI technology encompasses change detection and change captioning, each with its limitations in providing comprehensive interpretation. To address this, we propose an interactive Change-Agent which integrates a multi-level change interpretation (MCI) model as eyes and a large language model (LLM) as the brain. Our Change-Agent can follow user instructions to achieve comprehensive change interpretation and insightful analysis according to user instructions, such as change detection and change captioning, change object counting, change cause analysis, etc. Our proposed MCI model contains two branches of pixel-level change detection and semantic-level change captioning, in which multiple BI-temporal Iterative Interaction (BI3) layers utilize Local Perception Enhancement (LPE) and the Global Difference Fusion Attention (GDFA) modules to enhance the model’s discriminative feature representation capabilities. To train the MCI model, we build the LEVIR-MCI dataset with change masks and captions of bi-temporal images. Extensive experiments demonstrate the effectiveness of the proposed change interpretation model and highlight the promising potential of our Change-Agent in facilitating comprehensive and intelligent interpretation of surface changes. We will make our dataset and codebase of the change interpretation model and Change-Agent publicly available to facilitate future research at https://github.com/Chen-Yang-Liu/Change-Agent
Index Terms:
Interactive Change-Agent, change captioning, change detection, multi-level change interpretation, large language model.
I Introduction
The changes in the Earth’s surface, as a dynamic indicator of the Earth’s system, profoundly affect the planet’s evolution and the survival of humankind. Observing and analyzing these changes is pivotal for advancing sustainable human development [1]. Remote sensing (RS) satellite imagery, offering a unique “God’s perspective,” emerges as an effective tool for monitoring Earth’s dynamic changes. Remote sensing image change interpretation (RSICI) has emerged as a significant research focus [2]. RSICI aims to identify and analyze changes from images captured at different times in the same geographical area. It can provide decision-making support for environmental protection [3], urban planning [4], resource management [5], etc.
Current RSICI technology primarily encompasses change detection and change captioning. Change detection accurately localizes the spatial location of surface changes at the pixel level [6, 7, 8, 9, 10]. In contrast, change captioning aims at articulating the attributes and meanings of changes using natural language, emphasizing the semantic-level understanding of changes [11, 12, 13, 14, 15]. Despite significant advancements in both areas, their respective limitations prevent us from obtaining comprehensive change interpretation information by utilizing a single technology. Specifically, while change detection can accurately localize the changed area, it lacks a deeper understanding of changes’ underlying meanings, such as the characteristics of ground objects or the spatial relationships between them. Conversely, change captioning can provide rich semantic-level interpretation information but may fall short in providing precise pixel-level change localization. Therefore, there is a pressing need to explore a multi-level change interpretation (MCI) approach providing both pixel-level and semantic-level change information, facilitating precise change localization while delving into the essence and implications of the changes.
Moreover, practical applications often require comprehensive analysis and further processing of both pixel-level and semantic-level change interpretation results to meet specific needs. For instance, users may require statistical analysis of changed object numbers, consuming significant time and effort from researchers and demanding users’ technical proficiency.
To address these challenges, we propose an interactive Change-Agent based on a novel MCI model and a large language model (LLM). Fig.1 illustrates the advancements achieved by our Change-Agent compared to previous technologies. Different from the previous single technology, with the help of the MCI model, our Change-Agent can simultaneously achieve precise pixel-level change detection and semantic-level change captioning, thereby providing users with comprehensive change interpretation information. Notably, our Change-Agent boasts interactive capabilities, enabling users to communicate their queries or requirements regarding surface changes. Leveraging the MCI model and the LLM, the Change-Agent can adeptly interpret changes according to user needs, intelligently analyze and process the change interpretation information, and ultimately deliver tailored outcomes aligning with user expectations. The interactive Change-Agent bridges the gap between users and remote sensing expertise.
Figure 1: The comparison between previous single technology and our Change-Agent. Our Change-Agent can simultaneously achieve precise pixel-level change detection and semantic-level change captioning. Besides, it boasts interactive capabilities, enabling users to communicate their queries or requirements regarding surface changes.
As illustrated in Fig.1 (b), the MCI model is a pivotal cornerstone toward the realization of the Change-Agent. It enables the Change-Agent to perceive visual changes comprehensively, serving as the change perception foundation of the Change-Agent. However, there has been no prior work investigating the multi-task learning of change detection and change captioning. To address this gap, we built the LEVIR-MCI dataset. Each pair of bi-temporal remote sensing images in this dataset contains both pixel-level change detection masks and semantic-level change descriptions. Building upon this dataset, we propose a novel MCI model with two branches of change detection and change captioning. Within the two branches, we propose BI-temporal Iterative Interaction (BI3) layers utilizing Local Perception Enhancement (LPE) and the Global Difference Fusion Attention (GDFA) modules to enhance the model’s discriminative feature representation capabilities. The trained MCI model equips our Change-Agent with eyes, enabling the agent to realize comprehensive change interpretations of surface changes.
Recent advancements in LLMs such as ChatGPT [16] and Llama2 [17] have been remarkable. Trained on vast corpora, LLMs have acquired extensive knowledge and demonstrated powerful capabilities in instruction comprehension, planning, reasoning, question answering, and text generation [18, 19]. These notable progress have provided robust support for our LLM-based Change-Agent. In this work, the LLM serves as the brain of the Change-Agent, assuming a pivotal role in internal scheduling, planing, and user-intent understanding. Integration with LLMs imbues our agent with enhanced flexibility and intelligence, enabling a nuanced understanding of user intentions and provision of customized change interpretation and intelligent analysis services such as changed object counting, estimation of change causes, prediction of future changes, etc.
Our Change-Agent facilitates accurate and comprehensive change interpretation of RS images. It reduces the workload and time costs for researchers while enhancing the efficiency and convenience of interaction between users and remote sensing data. Extensive experiments demonstrate the effectiveness of the proposed MCI model and highlight the promising potential of our Change-Agent. Besides, the integration of the MCI model and LLM presents novel perspectives for analyzing surface changes.
Our contributions can be summarized as follows:
  • We built a multi-level change interpretation dataset. The dataset contains bi-temporal images as well as diverse change detection masks and descriptive sentences. It provides a crucial data foundation for exploring multi-task learning for change detection and change captioning.
  • We propose a dual-branch MCI model, which can provide pixel-level and semantic-level change interpretation. Besides, we propose BI3 layers with LPE and GDFA to enhance the change interpretation capability of the model. Experiments validate the effectiveness of our method.
  • Leveraging the MCI model and an LLM, we construct a Change-Agent that achieves an interactive and comprehensive interpretation and analysis of surface changes. It has intelligent dialogue and customized service capabilities, opening up new opportunities for intelligent remote sensing applications.
II Related work
A comprehensive change interpretation model should provide both pixel-level and semantic-level change information. However, previous work has primarily focused on either change detection or change captioning alone. Change detection aims to generate pixel-level change maps revealing the locations of changes in bi-temporal images, while change captioning aims to describe the changes via language. In this section, we will briefly review these approaches.
II-A Remote Sensing Change Detection
Early change detection approaches were traditionally categorized into algebra-based, transformation-based, and classification-based methods. These methods relied on techniques such as change vector analysis (CVA) [20, 21], principal component analysis (PCA) [22, 23], multivariate alteration detection (MAD) [24, 25], and post-classification comparisons [26, 27]. Traditional change detection methods have laid the foundation for modern techniques in remote sensing change detection [28].
With the emergence of deep learning, the field of remote sensing change detection has witnessed significant advancements [29, 30, 31, 32, 33]. Compared to traditional methods, deep learning significantly improves the performance of change detection with its superior multi-temporal feature learning capability. These methods utilize various architectures including convolutional neural networks (CNNs) [34, 35, 36, 37], recurrent neural networks (RNNs) [38, 39, 40, 41, 42], auto-encoders [43, 44, 45], and Transformers [46, 47, 6, 48, 49, 50]. For instance, Chen et al. [46] introduced a bi-temporal image Transformer (BIT) [46], where images are represented as semantic tokens, and the Transformer encoder and decoder refine visual tokens to identify changed areas efficiently and effectively. Li et al. [6] proposed an end-to-end TransUNetCD model based on Transformer and UNet structure. The model utilizes a Transformer encoder to extract global contextual features. A CNN-based decoder upsamples the multi-scale features to generate the change map. Bandara et al. proposed a pure transformer model named ChangeFormer [51], whose backbone is a siamese SegFormer [52]. Zhang et al. [9] propose a deep siamese network with multi-scale multi-attention. They enhance global representation using a contextual attention module and fuse multi-scale features from the siamese feature extractor for detecting objects of varying sizes and irregularities. Some surveys, such as [2, 1, 5] provide a more comprehensive review of change detection in deep learning.
II-B Remote Sensing Change Captioning
Remote sensing change captioning (RSCC) is a recently emerged multimodal task involving remote sensing image processing and natural language generation. Given the intricacy of this task, prevailing methodologies predominantly leverage deep learning techniques. Early methods in this domain largely drew inspiration from the encoder-decoder framework commonly utilized in image captioning tasks. Within this framework, the encoder is responsible for extracting features from bi-temporal images and performing the difference-aware fusion of these features. The decoder typically employs recurrent neural networks or Transformers to convert visual features into natural language. Compared to tasks like image captioning, which extract semantic representations from single images, change description faces the significant challenge of extracting robust semantic change representations.
Chouaf and Hoxha et al. [11, 12] pioneered remote sensing change captioning by introducing a model that employs a pre-trained CNNs [53] as the encoder and tried to utilize an RNN and Support Vector Machines (SVMs) as the decoder to generate textual descriptions. To improve the change-awareness of the model, they proposed an image-based fusion strategy and a feature-based fusion strategy. To facilitate research on change captioning, Liu et al. [13] introduced a large LEVIR-CC dataset and benchmarked some methods from other fields. Besides, they proposed a Transformer-based method named RSICCformer. The approach incorporates Siamese cross-encoding modules and multistage bitemporal fusion modules, emphasizing change regions through differencing features and capturing multiple changes of interest.
In a further advancement, some attention-based improvements have emerged [54, 14, 55]. For instance, PSNet [54] enhances the perception of changed objects of varying sizes using a progressive scale-aware network with difference perception layers and scale-aware reinforcement modules. Chang et al. [14] proposed an attention network containing a hierarchical self-attention module followed by a residual unit for identifying change-related attributes and constructing the semantic change embedding. Different from the encoder-decoder framework, Liu et al. [15] proposed a new paradigm to decouple the RSCC task by designing an image-level classifier and a feature-level encoder. Besides, a multi-prompt learning strategy is proposed to effectively exploit the potential of the LLM for captioning.
Figure 2: Examples of the LEVIR-MCI dataset. Each pair of bi-temporal images is provided with a change detection mask and one of the five sentences describing changes. In the change detection mask, changed buildings are highlighted in red, while changed roads are depicted in yellow.
III LEVIR-MCI Dataset
To cultivate a robust MCI model, foundational to the Change-Agent, we propose an MCI dataset named LEVIR-MCI (LEVIR Multi-level Change Interpretation)  [^1] LEVIR is our laboratory name: the Learning, Vision, and Remote Sensing Laboratory., featuring both pixel-level change information in the form of change detection masks and semantic-level insights encapsulated in descriptive sentences. The LEVIR-MCI dataset contains 10077 bi-temporal images. Each image has a spatial size of 256 × 256 pixels with a high resolution of 0.5 m/pixel and has a corresponding annotated mask and five annotated captions. The dataset will be publicly available at: https://github.com/Chen-Yang-Liu/Change-Agent
The proposed LEVIR-MCI dataset is an extension of our previously established change captioning dataset, LEVIR-CC [13]. In this dataset, we further provided each pair of bi-temporal images with additional change detection masks highlighting changed roads and changed buildings. Unlike its predecessor, the LEVIR-MCI dataset provides diverse annotations from different interpretation perspectives for each pair of images, further enhancing its utility for comprehensive change interpretation.
Fig. 2 presents some examples from our proposed LEVIR-MCI. Each pair of bi-temporal images within the dataset is annotated by a change detection mask delineating pixel-level alterations, along with one of five descriptive sentences elucidating the semantic-level changes. Such complementary annotations provide a multifaceted view of change interpretation, thereby bolstering the comprehensiveness of change interpretation models. To facilitate a nuanced understanding of the dataset, and considering our previous research has analysed the change captions, we will elucidate additional insights into the annotated change detection masks.
Figure 3: Distribution of the scale and deformation of changed roads and buildings. The ’area’ on the horizontal axis represents the area of a single object, and the ’bbox_area’ on the vertical axis represents the area of the corresponding rectangular bounding box. The dispersion of points offers insights into the diversity of object scale and deformation.
TABLE I: The Number of changed roads and buildings in the LEVIR-MCI dataset. The dataset comprises over 40,000 annotated instances of changed roads and buildings.
Set Number of changed objects Number of images with changed objects Average number of objects per image
Road Building Road Building Road Building
Training-set 3457 26155 2364 2972 1.46 8.80
Validation-set 611 4709 424 567 1.44 8.31
Test-set 934 8514 603 878 1.55 9.70
Total dataset 5002 39378 3391 4417 1.48 8.92
III-A Number of Changed Objects
In Table I, we meticulously analyse changed object masks in the LEVIR-MCI dataset. The dataset comprises over 40,000 annotated instances of changed roads and buildings. While the number of changed roads is fewer than buildings, generally exhibit longer spans and cover larger areas, as evidenced in Fig. 3. With such a vast collection of annotated masks, our LEVIR-MCI dataset can also serve as a fertile data ground for developing innovative multi-category change detection methodologies.
III-B Scale and Deformation of Changed Objects
An insightful examination of object scale and deformation is conducted through an analysis of road and building areas and the corresponding rectangular bounding boxes, as illustrated in Fig. 3. The ’area’ on the horizontal axis represents the area of a single object, and the ’bbox_area’ on the vertical axis represents the area of the corresponding bounding box. The dispersion of points offers insights into the diversity of object scale and deformation. Notably, points of roads exhibit a relatively broader spectrum with dispersed diversity, attributable to their narrow and curved nature. In contrast, points of buildings display a more concentrated distribution, reflecting their predominantly rectangular form. This analysis elucidates the intricate interplay between object scale and shape, shedding light on the nuanced characteristics of changed objects.
Figure 4: The overview of Change-Agent. The Change-Agent is equipped with an MCI model and an LLM, serving as its eyes and brain, respectively. The proposed LEVIR-MCI dataset provides a data foundation for training the MCI model.
In Fig. 4, we present an overview of the proposed Change-Agent. Comprising a MCI model as eyes and an LLM as the brain, the Change-Agent is endowed with visual perception and cognitive capabilities. This integration enables the agent to achieve an interactive and comprehensive interpretation and analysis of surface changes.
a comprehensive understanding and interpretation of surface changes, empowering the agent to provide insightful analyses and decision-making support. The MCI model facilitates robust and comprehensive change interpretation capabilities, while the LLM orchestrates the planning and information processing aspects of the agent. In the subsequent sections, we delve into the construction and training of the multi-level change interpretation model and elucidate the LLM-based scheduling of the Change-Agent.
IV METHODOLOGY
In Fig. 4, we present the overview of the proposed Change-Agent. We utilize the MCI model as eyes and an LLM as the brain to build our Change-Agent. The MCI model enables the Change-Agent to perceive visual changes comprehensively, serving as the change perception foundation of the Change-Agent. The constructed LEVIR-MCI dataset supports the multi-task training of the MCI model, which contains a change detection branch and a change captioning branch. As another pivotal role of the agent, the LLM can use its inherent rich knowledge to implement agent scheduling and provide insightful analyses and decision-making support. In the subsequent section, we will introduce the composition of the MCI model and the LLM-based scheduling of the agent.
IV-A Multi-level Change Interpretation Model
The MCI model serves as a central component of the Change-Agent, tasked with extracting and interpreting change information from bi-temporal remote sensing images. The proposed LEVIR-MCI dataset lays the groundwork for training the MCI model. The MCI model adopts a dual-branch architecture with a shared bottom, focusing on two pivotal tasks: change detection and change captioning. Specifically, a Siamese backbone network extracts multi-scale visual features from bi-temporal images, enabling learning in two branches. The lower-level features offer detailed information, while higher-level features are rich in semantics. Our change detection branch utilizes multi-scale features to refine change mask predictions, while the change captioning branch leverages the highest-level visual features to generate descriptive sentences. Through multi-task learning, we train a robust MCI model capable of simultaneously generating change detection masks and change captions.
IV-B Bi-temporal Iterative Interaction Layer
In the change detection and change captioning branches, we propose a novel BI-temporal Iterative Interaction (BI3) layer to effectively enhance and fuse bi-temporal features. The structure of the BI3 layer is illustrated in Fig. 5. The BI3 layer utilizes the Local Perception Enhancement (LPE) module and the Global Difference Fusion Attention (GDFA) module to extract the discriminative features of interest.
Figure 5: The structure of the Bi-temporal Iterative Interaction (BI3) layer. The BI3 layer utilizes the Local Perception Enhancement (LPE) and the Global Difference Fusion Attention (GDFA) to enhance the features of interest.
Motivated by the analysis presented in Section III-B regarding the various scales and deformations of roads and buildings, we propose the LPE module with a residual structure. The LPE module employs convolution kernels of different sizes to extract multiple feature maps across different scales. This design enriches the diversity of feature information and improves the model’s local feature perception capabilities. Assuming the bi-temporal features are denoted as ${X_{T_{1}}}\in\mathbb{R}^{N\times C}$ and ${X_{T_{2}}}\in\mathbb{R}^{N\times C}$, the LPE module can be formulated as follows:
$\mathrm{\Phi_{LPE}}(X_{T_{i}})$ $=\mathrm{ReLU(BN(conv_{1\times 1}}(F)))+{X_{T_{i}}}$ (1)
${F}$ $=\mathrm{concat}([{X_{s_{1}}},{X_{s_{2}}},{X_{s_{3}}}])$ (2)
${X_{s_{1}}}$ $=\mathrm{conv}_{3\times 3}({X_{T_{i}}})$ (3)
${X_{s_{2}}}$ $=\mathrm{conv}_{5\times 1}({X_{T_{i}}})$ (4)
${X_{s_{3}}}$ $=\mathrm{conv}_{1\times 5}({X_{T_{i}}})$ (5)
where $\mathrm{conv}_{1\times 1}$, $\mathrm{conv}_{3\times 3}$, $\mathrm{conv}_{5\times 1}$, $\mathrm{conv}_{1\times 5}$ represent $1\times 1$, $3\times 3$, $5\times 1$, and $1\times 5$ convolution kernels. concat stands for concatenation operation, BN (Batch Normalization) [56] for normalization, and ReLU (Rectified Linear Unit) for nonlinear activation function.
The GDFA module leverages differencing features to generate spatial attention weights and perform the interaction and fusion between features. This facilitates the model to focus on changes of interest and ignore irrelevant disturbances. Taking the left GDFA module in Figure 5 as an example, assuming $X_{lp_{1}}\in\mathbb{R}^{N\times C}$ and $X_{lp_{2}}\in\mathbb{R}^{N\times C}$ are the bi-temporal features obtained from the LPE module, the GDFA module can be formulated as follows:
$\mathrm{\Phi_{GDFA}}(X_{lp_{1}},X_{lp_{2}})$ $=\sigma({\frac{{Q}{K}^{T}}{\sqrt{d}}})V$ (6)
$Q$ $=X_{lp_{1}}W_{q}+b_{q}$ (7)
$K$ $=GW_{k}+b_{k}$ (8)
$V$ $=GW_{v}+b_{v}$ (9)
$Di$ $=X_{lp_{2}}-X_{lp_{1}}$ (10)
$G=\mathrm{\Phi_{DTF}}(Dif,$ $X_{L_{1}})=(Di*X_{L_{1}})W_{d}+b_{d}$ (11)
where $\sigma$ is the softmax function, $W_{q},W_{k},W_{v},W_{d}\in\mathbb{R}^{C\times d}$ are trainable weight matrices. $b_{q},b_{k},b_{v},b_{d}\in\mathbb{R}^{d}$ are trainable bias. $C$ denotes the dimension of $X_{LP_{i}}$ (i=1,2). $d$ is the dimension of transformed features.
Through the combination of LPE module and GDFA module, the BI3 layer improves the model’s feature representation and change discrimination capabilities. Following the LPE and GDFA modules, Layer Normalization (LN) [57] is applied to the bi-temporal features. Subsequently, a Multi-Layer Perceptron (MLP) with residuals further refines the normalized features to obtain enhanced bi-temporal features.
Figure 6: The structure of the change detection branch. “DeConv” denotes the deconvolution. The multi-scale features are gradually fused from bottom to top for refined mask prediction.
Figure 7: The structure of the change captioning branch. The domain bridging module further bridges the conversion from the visual domain to the textual domain. The Transformer decoder generates captions.
IV-C Change Detection and Captioning Branch
The structure of the change detection branch is depicted in Fig. 6. Leveraging multi-scale bi-temporal features extracted from the backbone network, this branch facilitates refined mask prediction. Specifically, multiple Bi-temporal Iterative Interaction (BI3) layers with residual connections iteratively enhance and refine bi-temporal high-level features, effectively capturing semantic changes. Since the low-level features contain more detailed information, they are crucial for refined change boundary detection. We further incorporate multiple Convolution-Based Bi-temporal Fusion (CBF) modules for bi-temporal feature fusion across four scales. Subsequently, through deconvolution, features are progressively integrated from bottom to top, bolstering the model’s discrimination capability for changes and enhancing change detection accuracy. For the input $X_{T_{1}}^{s},X_{T_{2}}^{s}(s=1,2,3,4)$, the CBF module can be represented as follows:
$\mathrm{\Phi_{CBF}}(X_{T_{1}}^{s},X_{T_{2}}^{s})$ $=\mathrm{conv_{1\times 1}(ReLU(BN(conv_{3\times 3}}(F))))$ (12)
$F$ $=\mathrm{concat}([X_{T_{1}}^{s},S,X_{T_{2}}^{s}])$ (13)
$S$ $=\mathrm{conv}_{3\times 3}(X_{T_{2}}^{s}-X_{T_{1}}^{s})+\mathrm{cos}(X_{T_{1}}^{s},X_{T_{2}}^{s})$ (14)
where $\mathrm{cos}(\cdot)$ denotes cosine similarity operation.
The change captioning branch focuses on semantic-level change interpretation, translating visual changes into textual descriptions, as depicted in Fig. 7. This branch leverages multiple BI3 layers to interactively process high-level semantic features extracted from the backbone network, thereby obtaining bi-temporal visual features revealing changes of interest. Subsequently, a convolution-based domain bridging module further processes the bi-temporal features to facilitate the transition from the visual domain to the textual domain. Finally, these processed features are fed into a Transformer decoder to generate descriptive sentences elucidating the changes. The processing flow of the domain bridging module can be expressed as follows:
$\mathrm{\Phi_{DB}}(X_{T_{1}},X_{T_{2}}))=F_{1}+\mathrm{conv_{1\times 1}}(F_{2})$ (15)
${F_{1}}$ $=\mathrm{conv_{1\times 1}(concat}([X_{T_{1}},X_{T_{2}})])$ (16)
$F_{2}$ $=\mathrm{conv_{3\times 3}(ReLU(BN(conv_{1\times 1}}(F_{1}))))$ (17)
Figure 8: The LLM plans how to complete the task according to the user’s instructions. We provide a suite of Python tools. The LLM drafts Python programs, subsequently executed by a Python interpreter to accomplish tasks beyond its inherent capabilities.
IV-D LLM: Brain of Change-Agent
Recently, LLMs have garnered significant attention for their remarkable achievements. By harnessing extensive web knowledge, LLMs have demonstrated substantial promise in attaining human-level intelligence, exhibiting noteworthy capabilities in tasks encompassing instruction comprehension, planning, reasoning, and proficient natural language interactions with humans [58, 59, 60, 61]. Building upon these advancements, there has been a growing research field that employs LLMs as central controllers for constructing autonomous agents [62, 63, 64, 65].
Inspired by these developments, we harness the capabilities of LLMs as the cognitive core responsible for orchestrating the scheduling of our Change-Agent. As depicted in Fig. 8, the LLM meticulously plans task execution based on user instructions. Despite excelling at text-related tasks, LLMs lack inherent visual perception capabilities. To bridge this gap and enable change interpretation and analysis akin to human capability, we equip our agent with a suite of Python tools. These tools include a visual feature extraction backbone, change detection branch, change captioning branch, and relevant Python libraries. Leveraging these tools, the LLM autonomously crafts Python programs, subsequently executed by a Python interpreter to accomplish tasks beyond its intrinsic capabilities. Subsequently, the LLM processes the results and furnishes feedback to the user. In addition to furnishing change masks and descriptions, our agent can implement object counting, image processing, change cause estimation, prediction of future changes, and more. In essence, our Change-Agent adeptly provides users with pertinent information, harnessing the rich knowledge of the LLM in tandem with the provided tools.
To facilitate the accurate generation of formatted Python code and tool invocation by the LLM, a meticulously crafted text prompt is essential. Inspired by the prompt utilized in [63], we devised a text prompt instructing the model on proper tool utilization. This prompt is input to the LLM in the role of a system instruction, as shown in Fig. 9.
Figure 9: The prompt input to LLM in the role of a system instruction.
V EXPERIMENTS
V-A Objective Function for Multi-task Training
During the training phase of our MIC model, we perform multi-task learning on both change detection and change captioning. Utilizing the meticulously constructed LEVIR-MCI dataset, we iteratively update the model parameters in a supervised learning framework.
For the change detection, we utilize the cross-entropy loss function to quantify the dissimilarity between predicted change masks and ground truth annotations. Specifically, the loss function can be mathematically defined as follows:
$\mathcal{L}_{\text{det}}=-\frac{1}{H\times W}\sum_{h=1,w=1}^{H,W}\sum_{j=1}^{C}y_{hw}^{(j)}\log(p_{hw}^{(j)})$ (18)
where $C$ is the number of classes. $H$ and $W$ are the height and width of the change mask. $\bm{p}_{hw}$ = [ $p_{hw}^{(0)},...,p_{cls}^{(C)}$] is the predicted probability vector at the pixel position (h,w)). $\bm{\widetilde{y}}_{hw}$ = [ $\widetilde{y}_{hw}^{(0)},...,\widetilde{y}_{hw}^{(C)}$] is the one-hot vector representation of the ground-truth at the pixel position (h,w)).
Similarly, for change captioning, we utilize the cross-entropy function to compute the loss between predicted sentences and their corresponding ground truth sentences. The formulation of the captioning loss function can be expressed as follows:
$\mathcal{L}_{\text{cap}}=-\frac{1}{L}\sum_{l=1}^{L}\sum_{v=1}^{V}y_{l}^{(v)}\log(p_{l}^{(v)})$ (19)
where $L$ represents the total number of word tokens, $V$ denotes the vocabulary size, $y_{l}^{(v)}$ denotes the ground truth label indicating whether $l$-th word belongs to vocabulary index $v$, and $p_{l}^{(v)}$ denotes the predicted probability of $l$-th word being assigned to vocabulary index $v$.
To balance the losses of these two tasks during model training, we adopt a normalization approach, scaling the losses of both tasks to the same order of magnitude. This ensures that each task contributes equally to the overall loss function and facilitates effective simultaneous optimization of the change detection and change captioning. Mathematically, the total loss can be expressed as follows:
$\mathcal{L}_{total}=\frac{\mathcal{L}_{det}}{\mathrm{detach}(\mathcal{L}_{det})}+\frac{\mathcal{L}_{cap}}{\mathrm{detach}(\mathcal{L}_{cap})}$ (20)
where detach($\cdot$) denotes the operation to detach the gradient flow of the loss.
V-B Evaluation Metrics
To evaluate the performance of our MCI model in detecting multi-category changes in remote sensing images, we use the Mean Intersection over Union (MIoU). The MIoU metric assesses the spatial overlap between predicted change masks and ground-truth masks, providing a reliable measure of pixel-level change detection accuracy. Mathematically, the calculation of MIoU is defined as:
$\text{MIoU}=\frac{1}{\text{N}}\sum_{i=1}^{\text{N}}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}$ (21)
Where N is the total number of classes. For the $i$-th class, $\text{TP}_{i}$, $\text{FP}_{i}$, and $\text{FN}_{i}$ represent the number of true positive, false positive, and false negative, respectively.
To evaluate the performance of our MCI model in describing changes between bi-temporal images, we adopted several evaluation metrics commonly used in previous change captioning studies [66, 14, 13, 12]. These metrics are as follows:
  • BLEU-n [67]: The BLEU-n metric measures the n-gram similarity between generated sentences and ground-truth sentences. We adopt n=1, 2, 3, and 4 to compute the BLEU-n score.
  • METEOR [68]: The METEOR computes the harmonic mean of precision and recall of single words. Besides, it incorporates a penalty factor to consider the fluency of generated sentences.
  • ROUGEL [69]: The ROUGEL metric measures the similarity of the longest common subsequence between generated sentences and ground-truth sentences. It is more concerned with the recall rate and is suitable for evaluating long sentences.
  • CIDEr-D [70]: The CIDEr metric treats each sentence as a document and represents it in the form of Term Frequency Inverse Document Frequency (TF-IDF) vectors. After computing TF-IDF for each n-gram, CIDEr-D is obtained by performing the cosine similarity.
By employing this comprehensive suite of evaluation metrics, we aim to provide a thorough assessment of the change detection and captioning capabilities of our MCI model. Higher scores across these metrics indicate more accurate change masks or superior sentence quality.
V-C Experimental Details
We implemented all models using the PyTorch deep learning framework and trained them on the NVIDIA GTX 4090 GPU. During training, we minimized the loss function defined by Equation 20 and employed the Adam optimizer [71] with an initial learning rate of 0.0001 to optimize the model parameters. We set the maximum number of epochs to 200 and utilized word embeddings with a dimensionality of 512. The backbone for feature extraction is the Siamese weight-shared Segformer-B1 [52]. Initially, we trained the backbone until the sum of BLEU and MIoU indicators did not increase for 50 consecutive epochs. Subsequently, we froze the backbone network and continued training the two branches separately.
TABLE II: The comprehensive performance comparisons on the LEVIR-MCI dataset. There is a lack of methods that simultaneously address change detection and change captioning. To assess the efficacy of our model in handling both tasks, we compared our method with established methods in each respective field.
Method MIoU BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D
Change Detection FC-EF [35] 84.54
FC-Siam-Conc [35] 85.42
FC-Siam-Di [35] 85.20
BIT [46] 84.45
Change Captioning Capt-Rep-Diff [66] 72.90 61.98 53.62 47.41 34.47 65.64 110.57
Capt-Att [66] 77.64 67.40 59.24 53.15 36.58 69.73 121.22
Capt-Dual-Att [66] 79.51 70.57 63.23 57.46 36.56 70.69 124.42
DUDA [66] 81.44 72.22 64.24 57.79 37.15 71.04 124.32
MCCFormer-S [72] 79.90 70.26 62.68 56.68 36.17 69.46 120.39
MCCFormer-D [72] 80.42 70.87 62.86 56.38 37.29 70.32 124.44
RSICCFormerc [13] 83.09 74.32 66.66 60.44 38.76 72.63 130.00
PSNet [54] 83.86 75.13 67.89 62.11 38.80 73.60 132.62
Chg2Cap [14] 86.14 78.08 70.66 64.39 40.03 75.12 136.61
MCI Ours 86.43 86.68 78.75 71.74 65.95 40.80 75.96 140.29
V-D Multi-level Change Interpretation Performance
Currently, there is a dearth of methods that simultaneously address change detection and change captioning. To assess the efficacy of our model in handling both tasks, we conducted comprehensive performance evaluations by comparing against established methods in each respective field on the proposed dataset.
For the change detection evaluation, we have benchmarked several well-known methods, including FC-EF [35], FC-Siam-Conc [35], FC-Siam-Di [35], and BIT [46]. Meanwhile, for the change captioning assessment, we compare our model with several existing state-of-the-art (SOTA) methods, including Capt-Rep-Diff [66], Capt-Att [66], Capt-Dual-Att [66], DUDA [66], MCCFormer-S [72], MCCFormer-D [72], RSICCFormer [13] and Chg2Cap [14].
The comprehensive performance comparisons, as shown in Table II, underscore the superiority of our proposed methodology across both change detection and change captioning tasks. Our method not only achieves simultaneous proficiency in both tasks but also surpasses previous methods. Specifically, our model exhibits superior performance in change detection with a +1.01% improvement on MIoU. Moreover, in terms of change captioning, our model exhibits a noteworthy advancement, showcasing a +1.56% improvement on BLEU-4 and a substantial +3.68% improvement on CIDEr-D.
Through the above performance analysis, we gain insights into the model’s ability to generate pixel-level and semantic-level interpretation information of changes observed in remote sensing images. This highlights the efficacy of our proposed MCI model in providing a comprehensive and in-depth analysis of surface changes.
TABLE III: Ablation studies on the effectiveness of the proposed BI3 layer, where the bolded results are the best.
Method MIoU BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D
Baseline 86.40 84.34 75.83 68.72 63.05 39.22 73.82 133.58
  +GDFA 86.41 85.32 77.45 70.63 65.10 39.59 74.45 135.94
    +GDFA+LPE 86.43 86.68 78.75 71.74 65.95 40.80 75.96 140.29
Figure 10: The loss curve shows the imbalance in the training process of the two tasks when we add the detection loss and caption loss to form the total loss.
TABLE IV: Effect on the loss balancing strategy detailed in Equation 20. “W.o. balance” means that we simply combined the detection loss and caption loss to form the total loss to optimize the multi-task model parameters.
Method MIoU BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D
W.o. balence 85.28 86.81 78.95 71.84 65.87 40.61 75.92 139.75
W. balence 86.43 86.68 78.75 71.74 65.95 40.80 75.96 140.29
TABLE V: Comparative analysis of single-task and multi-task learning. While multi-task learning performs poorer change detection than single change detection training, it contributes to improving change captioning capabilities, particularly on critical BLEU-4 and CIDEr-D metrics.
Method MIoU BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGEL CIDEr-D
Single Change Detection 86.54
Single Change Captioning 86.56 78.70 71.69 65.86 40.53 75.87 139.92
Multi-task Training 86.43 86.68 78.75 71.74 65.95 40.80 75.96 140.29
V-E Ablation Studies
V-E1 BI3 Layer
As a critical component of our change interpretation model, the BI3 layer incorporates LPE and GDFA modules. The LPE module enhances feature representation through local perception enhancement,improving the model’s sensitivity to multi-scale changes. The GDFA module effectively integrates the differential information, allowing the model to capture the change area more accurately. In Table III, we conducted ablation experiments on the effectiveness of the LPE and GDFA modules. The baseline utilizes Transformer encoding layers to process bi-temporal features. The experimental results show that integrating the LPE and GDFA modules leads to improved performance in both change detection and change captioning tasks.
V-E2 Balancing Change Detection and Captioning
In our exploration of multi-task learning for change detection and change captioning, achieving an effective balance between these tasks posed a significant challenge. Initially, when we simply combined the individual losses $\mathcal{L}{det}$ and $\mathcal{L}{cap}$ to form the total loss and minimized it to optimize the multi-task model parameters, we observed large discrepancies in the magnitudes of these two losses, as illustrated in Fig. 10. This imbalance during training could potentially impact overall change interpretation performance. To address this issue, we adopt a balancing strategy detailed in Equation 20 to normalize the two losses and achieve the training balance between the two tasks. Table IV shows the impact of these two strategies on model performance. It’s evident that our balancing strategy plays a crucial role in harmonizing the training process, leading to more balanced performance across both change detection and captioning tasks.
In addition, we conducted a comparative analysis of single-task and multi-task learning, as shown in Table V. While multi-task learning yields poorer change detection performance compared to single change detection training, it contributes to improving change captioning capabilities, particularly on critical BLEU-4 and CIDEr-D metrics. This observation could be attributed to the shared knowledge and feature representation learning across tasks. Pixel-level change detection can be advantageous for accurately localizing changes when generating change descriptions. However, it’s worth noting that ground-truth change descriptions sometimes overlook minor changes, which could be detrimental to pixel-level change detection. The trade-off between change detection and change captioning underscores the complex interplay between pixel-level and semantic-level interpretation of surface changes.
In conclusion, achieving a balance between change detection and captioning, and fostering collaborative improvement between them, remains a subject worthy of further exploration. We look forward to discovering more effective methods to address this challenge in future research.
Refer to captionRefer to caption
Figure 11: The qualitative comparison results between the baseline and our model. For captioning results, we provide one of the five ground-truth reference sentences. Besides, more accurate and detailed words are marked in green, while red words are not. The results demonstrate that our model exhibits strong change interpretation capabilities in change detection and change captioning.
V-F Qualitative Change Interpretation Results
To illustrate the effectiveness of our proposed model in simultaneously generating change masks and change descriptions, Fig. 11 presents comparative results between our proposed model and the baseline model. The experimental results demonstrate the superior capabilities of our model in both change detection and captioning.
In terms of change detection, our model exhibits remarkable proficiency in identifying subtle changes in small buildings, as depicted in the first row of Fig. 11. Furthermore, our model can produce more refined road masks in rows 5 and 7 compared to the baseline. In contrast to the baseline, which erroneously detects some buildings in the sixth and eighth rows, our model performs admirably.
Regarding change captioning, in the fourth and eighth rows of Fig. 11, our model can accurately recognize and describe the changes in the bi-temporal images. Conversely, the baseline erroneously concludes that no changes have occurred, resulting in a critical error in the description. Furthermore, it is evident from the first four rows of Fig. 11 that our model can not only accurately identify the types of changing objects but also describe their spatial positions and the relationships between the changing objects and surrounding features.
Figure 12: Examples of conversations between a user and our Change-Agent. Our Change-Agent can understand users and carry out tasks such as change mask prediction, change captioning, change object counting, mask post-processing, change causes estimation, and future change prediction.
Figure 13: The executable Python function code generated by LLM when a user requests to only detect changed buildings and display it in red.
V-G Interaction Between User and Change-Agent
Leveraging the change interpretation model trained on the LEVIR-MIC dataset, we have constructed an LLM-based Change-Agent. The change interpretation model provides our agent with powerful change interpretation capability. The LLM enables our agents to understand user instructions and call tools to achieve interactive and comprehensive change interpretation and processing.
In Fig. 12, we illustrate two examples of conversations between a user and our Change-Agent. In the first conversation, our agent showcases its adeptness in understanding user requirements and subsequently carries out tasks such as change mask prediction, change captioning, change object counting, and mask post-processing. This highly automated processing greatly enhances user efficiency and provides users with a more convenient and intuitive experience in change interpretation. In the second conversation, our agent can describe the changes in the images with the help of change interpretation model, and use the rich knowledge inherent in the LLM to answer the user’s questions about the cause of the changes and the prediction of future changes. This will assist user decision-making.
Fig. 13 presents a specific example to illustrate the practical application value of our Change-Agent. When a user requests our agent to only detect changed building masks and display building areas in red, the LLM within our agent can accurately understand this instruction and generate executable Python function code.
Through these interactive capabilities, our Change-Agent serves as a versatile tool for facilitating nuanced and insightful analysis of changes within remote sensing images, fostering seamless communication and collaboration between users and remote sensing expertise.
VI Discussion and Future
Our research introduces the LEVIR-MIC dataset that lays a valuable data foundation for facilitating multi-task learning of change detection and change captioning and exploring the interconnection between two tasks. Besides, our Change-Agent opens up new avenues for change interpretation and brought a new thinking direction to the remote sensing field. However, there are still some issues worthy of in-depth exploration. Future work can be carried out from the following aspects:
  • Multi-Task Learning. Based on the proposed dataset, we initially explored multi-task learning of change detection and change captioning and found some challenges, such as the balance of two tasks. The intrinsic relationship between the two tasks still needs to be deeply explored. By designing better model structures and training strategies, future work can focus on further promoting the mutual promotion between the two tasks, thereby improving the overall interpretation performance.
  • Scheduling of Change-Agent. The system prompts designed by humans and inputted into the LLM play a crucial role in facilitating accurate agent scheduling. By optimizing the prompt design, we can enable the agent to more accurately capture the user’s intentions and plan a more reasonable and efficient task execution pathway.
  • Tool Expansion. Beyond change interpretation, we can provide the agent with additional models to enhance its capabilities to handle more complex and comprehensive remote sensing image processing tasks, thereby providing users with more comprehensive and professional interpretation services. Additionally, the updateability of tools in the agent allows us to introduce new tools and models, ensuring that our agent remain up-to-date.
  • Multi-Agent Systems. By introducing multiple agents with diverse capabilities and fostering collaboration among them, researchers can achieve more flexible and collaborative image interpretation work. It is helpful for building a more efficient and intelligent remote sensing system.
VII Conclusion
In this paper, we have addressed the critical need for comprehensive and intelligent interpretation of surface changes on Earth through the development of a novel Change-Agent. Our agent is propelled by a multi-level change interpretation model and an LLM. The change interpretation model endows the agent with robust visual change perception capabilities, while the LLM orchestrates the agent’s scheduling. The proposed change interpretation model encompasses dual branches of change detection and captioning, capable of providing both pixel-level change masks and semantic-level understanding in textual form. Within the dual-branch framework, we introduce a BI3 layer with LPE and GDFA to enhance the model’s discriminative feature representation capabilities. Additionally, we build a dataset comprising diverse change detection masks and descriptions to support model training. Experiments validate the efficacy of the proposed change interpretation model and highlight the promising potential of our Change-Agent in facilitating comprehensive and intelligent interpretation of surface changes.